Can someone help me solve this problem?
I need to solve the system by addition:
5x - 3y = 13
4x - 3y = 11
Thank you,
Sher

Found 2 solutions by funmath, luanaoh:
Answer by funmath(2933)

Hi Sher,
Solve the system by addition:
E1: 5x - 3y = 13
E2: 4x - 3y = 11
:
Multiply E2 by -1 and add E1+E2 and the y's will be eliminated so that you can solve for x.
:
E2: -1(4x-3y)=-1(11)----> -4x+3y=-11
:
5x-3y=13
-4x+3y=-11
__________
x+0=2
x=2
Substitute x=2 into either equation and solve for y.
5(2)-3y=13
10-3y=13
-10+10-3y=-10+13
-3y=3
-3y/-3=3/-3
y=-1
The solution is: (x,y)=(2,-1)
Happy Calculating, Sher!!!


Answer by luanaoh(1)

umm...i'm not sure if the answer would be right but..i will give it a try.
5x - 3y = 13
4x - 3y = 11
-------------
9x-6y=24 ...um...lets find y.
-9x
--------
-6y=-9x+24
now u devide by -6.
y=1 1/3x +24.

you plug that in to the y in any equation and if u get the same answer to the problem. it is a solution...
i think.